# A Versatile Pathology Co-pilot via Reasoning Enhanced Multimodal Large Language Model

### v1

@@ Abstract
Multimodal large language models (MLLMs) have emerged as powerful tools for computational pathology, offering unprecedented opportunities to integrate pathological images with language context for comprehensive diagnostic analysis. These models hold particular promise for automating complex tasks that traditionally require expert interpretation of pathologists. However, current MLLM approaches in pathology demonstrate significantly constrained reasoning capabilities, primarily due to their reliance on expensive chain-of-thought annotations. Additionally, existing methods remain limited to simplex application of visual question answering (VQA) at region-of-interest (ROI) level, failing to address the full spectrum of diagnostic needs such as ROI classification, detection, segmentation, whole-slide-image (WSI) classification and VQA in clinical practice. In this study, we present SmartPath-R1, a versatile MLLM capable of simultaneously addressing both ROI-level and WSI-level tasks while demonstrating robust pathological reasoning capability. Our framework combines scale-dependent supervised fine-tuning and task-aware reinforcement fine-tuning, which circumvents the requirement for chain-of-thought supervision by leveraging the intrinsic knowledge within MLLM. Furthermore, SmartPath-R1 integrates multiscale and multitask analysis through a mixture-of-experts mechanism, enabling dynamic processing for diverse tasks. We curate a large-scale dataset comprising 2.3M ROI samples and 188K WSI samples for training and evaluation. Extensive experiments across 72 tasks validate the effectiveness and superiority of the proposed approach. This work represents a significant step toward developing versatile, reasoning-enhanced AI systems for precision pathology.
@@ 1 Introduction
Figure 1: Overview of the SmartPath-R1. a. SmartPath-R1 is a versatile pathology co-pilot that can simultaneously addresses both ROI-level and WSI-level tasks across various anatomical regions while demonstrating robust pathological reasoning capability. b. We curate a dataset comprising 2,347,453 ROI-level and 188,154 WSI-level samples across 72 diverse tasks, enabling comprehensive model training and evaluation. c. Performance evaluation of MLLMs across 46 ROI-level tasks and 26 WSI-level tasks. * represents external validation tasks. SmartPath-R1 demonstrates exceptional performance over state-of-the-art MLLMs. d. The framework, comprising scale-dependent supervised fine-tuning and task-aware reinforcement fine-tuning, enables multimodal pathological reasoning within <think></think> tags. ROIs and WSIs are encoded into variable-length visual token sequences. Mixture-of-expert mechanism is used to address a spectrum of tasks, from fine-grained classification to WSI-level VQA.
@@
The emergence of multimodal large language models (MLLMs) [20, 21, 22, 23] has created new opportunities, yet two fundamental limitations persist. First, existing models often fail to achieve clinically meaningful reasoning. This is largely due to their reliance on supervised fine-tuning  [24, 22, 23], which optimizes the accuracy of prediction but fails to capture the sequential decision-making process inherent in pathological diagnosis, where experts progressively integrate visual cues and knowledge. As a result, models may learn superficial correlations rather than the critical reasoning pathways needed for complex cases. Second, existing approaches exhibit limited versatility in handling various computational pathology tasks. Although current MLLMs demonstrate preliminary capabilities in ROI-level VQA [22, 23], their capacity to concurrently process multiscale queries (ROI and WSI level) remains underdeveloped. These methods lack unified frameworks to integrate fundamental pathological tasks, including fine-grained classification, detection, and segmentation, with diagnostic reasoning. This fragmentation hinders the development of comprehensive diagnostic assistants capable of mirroring the multi-task analytical workflows of pathologists, where visual interpretation and semantic reasoning are intrinsically coupled across spatial hierarchies.
@@
To overcome these limitations, we present SmartPath-R1 (Figure 1), a MLLM trained on extensive datasets comprising 1,964,121 ROI samples and 179,897 WSI samples, with rigorous evaluation performed on 83,332 ROI and 8,257 WSI test samples. Our approach uniquely embeds diagnostic reasoning by formulating pathology interpretation as a sequential policy optimization process, where the model learns optimal evidence-gathering policies through reward signals mimicking pathologist decision trajectories. Crucially, our approach eliminates the need for expensive chain-of-thought supervision by mining the knowledge inherent in MLLMs, thereby enhancing scalability. Moreover, SmartPath-R1 unifies multiscale and multitask analysis via mixture-of-expert mechanism that dynamically handles ROI-level morphometrics to WSI-level topological features for different tasks.
@@ 2 Results
We evaluate various MLLMs across 72 tasks, encompassing 8 ROI-level classification tasks, 14 ROI-level pathological detection tasks, 14 ROI-level pathological segmentation tasks, 10 ROI-level VQA tasks, 9 WSI-level classification tasks, and 17 WSI-level VQA tasks.
@@
Pathological ROI-level classification is a fundamental task in computational pathology, aiming to automatically categorize pathology ROI images into predefined classes. Unlike generic image classification, this task demands strong reasoning capabilities to decipher subtle morphological patterns (e.g., nuclear pleomorphism, tissue architecture) and integrate domain-specific knowledge (e.g., grading criteria, staining artifacts). While traditional deep learning approaches excel at pattern recognition, integrating explicit reasoning mechanisms is increasingly vital to address challenges like label ambiguity, stain variability, and rare class generalization, ultimately bridging the gap between AI and clinical trustworthiness.
@@
In our comprehensive evaluation, SmartPath-R1 is benchmarked against four state-of-the-art MLLMs: Qwen2.5VL [25] (general-purpose vision-language model), LLaVA-Med [26] (biomedical-focused MLLM), Quilt-LLaVA [23] and PathoR1 [27] (specialized for pathology). As shown in Figure 2a-k, the evaluation of MLLMs across 8 distinct tasks, including 2 external validation datasets, reveals important insights into their capabilities. Qwen2.5VL surprisingly outperforms several domain-specialized models across multiple pathological datasets. This finding can be attributed to its extensive pretraining on diverse vision-language tasks, which fosters more robust feature representations capable of generalizing to unseen pathological patterns. PathoR1’s relatively limited performance highlights the challenges of overspecialization in pathological AI systems. While designed specifically for ROI-level VQA, its constrained training scope and modest sample size render it less adaptable to broader diagnostic tasks. In contrast, SmartPath-R1 achieves particularly strong results, where it outperforms the second-best approach Qwen2.5VL by 38.3% in average accuracy. This superior performance can be attributed to the models’ ability to capture tissue patterns and contextual relationships in these complex classification tasks.
@@
We present an example of classification task in Figure 2l. General-purpose Qwen2.5VL struggles with domain-specific reasoning (e.g., grading dysplasia), as it prioritize broad visual-language alignment over histopathologic logic. LLaVA-Med’s correct but terse output (“shows a tumor”) lackes explanatory depth, limiting its utility in clinical settings where justification is paramount. Quilt-LLaVA and PathoR1’s misclassification as (G) and (B) likely stem from isolated feature recognition without integrating stromal or nuclear atypia cues. SmartPath-R1 identifies the tissue as (J) Tumor, reasoning that the glandular morphology indicates advanced gastric adenocarcinoma, excluding early-stage or benign features (e.g., ulceration or regression tissue). This aligns with the clinical context of adenocarcinoma, where malignant glandular presence is a hallmark.
@@
Figure 2: Performance of MLLMs on ROI-Level Classification Tasks. a-c. Average performance of MLLMs based on accuracy (ACC) across 8 ROI-level classification tasks. d-k. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test [28]. Additional results are shown in Extended Data Table 3 and Tables 4-11. l. An example of ROI-level classification along with the results generated by various MLLMs.
@@
Figure 3: Performance of MLLMs on ROI-Level Detection Tasks. a-f. Average performance of MLLMs based on average precision at IoU threshold 0.5 (AP50) across 14 ROI-level detection tasks. g-t. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 13, Tables 14-27, Figure 8 and Figure 9. u. An example of ROI-level detection task along with the results generated by various MLLMs. Red rectangles indicate ground truth annotations, while predicted bounding boxes of each method are shown in their corresponding colors. The AP50 scores for each method are displayed in the lower-right corner, quantitatively demonstrating their detection accuracy relative to the ground truth.
@@
ROI-level detection involves the localization and classification pathological entities (e.g., cells, nuclei, tumor regions) in images. This task serves as the foundation for automated diagnosis, prognostic assessment, and biomedical research, enabling high-throughput analysis of morphological features that are often imperceptible to human observers. The clinical significance of this task necessitates not only detection accuracy but also interpretive reasoning—the ability to contextualize morphological patterns within pathological paradigms. For instance, distinguishing malignant nuclei from benign counterparts requires reasoning about spatial relationships (e.g., tumor infiltration), texture heterogeneity, and diagnostic criteria. Traditional deep learning approaches, while proficient in pattern recognition, often lack explicit reasoning mechanisms.
@@
Figure 3 presents a comparative evaluation of SmartPath-R1 with existing MLLMs and a state-of-the-art medical image detection model BiomedParse. The analysis leverages 14 benchmark datasets spanning nuclear, cell, and tissue detection, focusing on Average Precision at an IoU threshold of 0.5 (AP50) as the primary metric. The results demonstrate that SmartPath-R1 achieves a significantly higher average AP50 across all datasets (Figure 3a-f), indicating superior detection performance. Dataset-specific performances (Figures 3g-t) further reveal that SmartPath-R1 exhibits particularly strong performance on all the datasets (p < 0.001). This suggests that SmartPath-R1 possesses a greater capacity for reasoning about the complex visual cues necessary for accurate detection.
@@
The qualitative examples is shown in Figure 3u. SmartPath-R1’s bounding boxes tightly enclose the glands, demonstrating a precise understanding of gland boundaries and morphology. In contrast, other models often produce inaccurate or incomplete bounding boxes, indicating a failure to reason about the subtle visual cues that distinguish glands from surrounding tissue. These findings suggest that SmartPath-R1’s enhanced reasoning capabilities contribute to its superior performance on ROI-level detection tasks, enabling it to accurately localize objects even in challenging histological contexts.
@@
Figure 4: Performance of MLLMs on ROI-Level Segmentation Tasks. a-c. Average performance of MLLMs based on Dice score across 14 ROI-level segmentation tasks. d-q. Model performance on specific tasks. * represents external validation datasets. The violin plots visualize the performance distribution of different methods, where each violin’s width represents the data density. The blue horizontal line within each violin indicates the median value, while the white circle marks the mean point. Black vertical lines show the 95% CI. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 28 and Tables 29-30. r. An example of ROI-level segmentation task along with the results generated by various MLLMs. Red masks indicate ground truth annotations, while predicted masks of each method are shown in their corresponding colors. The Dice scores for each method are displayed in the lower-right corner, quantitatively demonstrating their segmentation accuracy relative to the ground truth.
@@
To assess the segmentation capabilities of SmartPath-R1, we conduct a comparative evaluation against existing models on 14 ROI-level segmentation datasets (Figure 4). The primary evaluation metric is the Dice score, which quantifies the overlap between predicted and ground truth segmentation masks. The results demonstrate that SmartPath-R1 achieves a significantly higher average Dice score across all datasets (Figures 4a-c) compared to other models, indicating superior segmentation performance. Analysis of dataset-specific performance (Figures 4d-q) reveals that SmartPath-R1 exhibits particularly strong performance on datasets such as BCSS, CoNIC2022, CRAG, DigestPath2019 Cell, DigestPath2019 Tissue, MIDOG Breast, MIDOG Neuroendocrine, Rings, SICAPv2, and WSS4LUAD (p < 0.001). This suggests that SmartPath-R1 possesses a greater capacity for reasoning about the complex visual cues necessary for accurate segmentation.
Figure 4r presents the qualitative examples. SmartPath-R1’s segmentation masks closely align with the lymphocyte boundaries, demonstrating a precise understanding of cell morphology. In contrast, other models often produce inaccurate or incomplete masks, indicating a failure to reason about the subtle visual cues that distinguish lymphocytes from surrounding tissue. These findings suggest that SmartPath-R1’s enhanced reasoning capabilities contribute to its superior performance on ROI-level segmentation tasks, enabling it to accurately delineate histological structures even in complex pathology image.
@@
Figure 5: Performance of MLLMs on ROI-Level VQA Tasks. a-b. Average performance of MLLMs based on accuracy (ACC) across 10 ROI VQA tasks. c-l. Model performance on specific tasks. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 31 and Table 32. m. An example of ROI-level VQA task along with the results generated by various MLLMs.
Pathological VQA is an advanced multimodal reasoning task that requires models to interpret medical images and generate accurate answers to free-form clinical questions. Unlike classification operates within a static label space (e.g., binary tumor grading or disease categorization) per dataset, VQA dynamically adapts to diverse diagnostic inquiries, thereby emulating the flexible and context-driven nature of real-world clinical decision-making. This task not only demands fine-grained visual-textual alignment but also produces more interpretable and clinically actionable outputs, making it suited for complex diagnostic scenarios where reasoning and justification are critical.
To benchmark the performance of SmartPath-R1 against existing MLLMs on fundamental ROI-level VQA task, we conduct a comparative evaluation on 10 datasets. Results demonstrate that SmartPath-R1 achieves a higher average accuracy (ACC = 0.683) compared to other models, including Qwen2.5VL (ACC = 0.406), LLaVA-Med (ACC = 0.246), Quilt-LLaVA (ACC = 0.307) and PathoR1 (ACC = 0.559) (Figures 5a-b). This suggests an advantage in general ROI-level understanding. Detailed analysis of task-specific performance (Figures 5c-l) reveals that SmartPath-R1 exhibits statistically significant (p < 0.05) improvements on tasks requiring more complex reasoning about histological structures.
In the illustrative ROI-level VQA example (Figure 5m), SmartPath-R1 correctly identifies the tissue type while other MLLMs fail, qualitatively demonstrating its enhanced reasoning capabilities. We attribute these gains to SmartPath-R1’s training on a large-scale dataset without procedural supervision, enabling the model to learn more robust features and develop more sophisticated reasoning skills.
@@
Figure 6 summarizes the accuracy of five MLLMs across 9 WSI classification tasks including 3 external ones. SmartPath-R1 consistently achieves the highest average accuracy (15.5 % absolute margin over the second-best method; p < 0.001, Wilcoxon signed-rank test), and demonstrates superior robustness on both internal (TCGA-BRCA, TCGA-COAD, TCGA-HNSC, TCGA-LGG, TCGA-LUAD, TCGA-LUSC) and external validation sets (BRACS, CAMELYON, CPTAC-NSCLC).
In the glioma subtyping example (Figure 6m ), competing models falter due to either insufficient pathological knowledge or perceptual errors in feature recognition, leading to misclassifications. SmartPath-R1 instead performs a deliberate, step-wise exclusion: absence of mitoses, microvascular proliferation and necrosis rules out Glioblastoma Multiforme; moderate hypercellularity and mild nuclear pleomorphism align with anaplastic astrocytoma. The model explicitly states “no evidence of… distinguishing it from glioblastoma multiforme”, yielding a coherent rationale that leads to the correct answer (B) Anaplastic Astrocytoma.
Figure 6: Performance of MLLMs on WSI-Level classification Tasks. a-c. Average performance of MLLMs based on accuracy (ACC) across 9 WSI-level classification tasks. d-l. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 34 and Table 35. m. An example of WSI-level classification task along with the results generated by various MLLMs.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have emerged as powerful tools for computational pathology, offering unprecedented opportunities to integrate pathological images with language context for comprehensive diagnostic analysis. These models hold particular promise for automating complex tasks that traditionally require expert interpretation of pathologists. However, current MLLM approaches in pathology demonstrate significantly constrained reasoning capabilities, primarily due to their reliance on expensive chain-of-thought annotations. Additionally, existing methods remain limited to simplex application of visual question answering (VQA) at the region-of-interest (ROI) level, failing to address the full spectrum of diagnostic needs such as ROI classification, detection, segmentation, whole-slide-image (WSI) classification and VQA in clinical practice. In this study, we present SmartPath-R1, a versatile MLLM capable of simultaneously addressing both ROI-level and WSI-level tasks while demonstrating robust pathological reasoning capability. Our framework combines scale-dependent supervised fine-tuning and task-aware reinforcement fine-tuning, which circumvents the requirement for chain-of-thought supervision by leveraging the intrinsic knowledge within MLLM. Furthermore, SmartPath-R1 integrates multiscale and multitask analysis through a mixture-of-experts mechanism, enabling dynamic processing for diverse tasks. We curate a large-scale dataset comprising 2.3M ROI samples and 188K WSI samples for training and evaluation. Extensive experiments across 72 tasks validate the effectiveness and superiority of the proposed approach. This work represents a significant step toward developing versatile, reasoning-enhanced AI systems for precision pathology.
@@ 1 Introduction
Figure 1: Overview of the SmartPath-R1. a. SmartPath-R1 is a versatile pathology co-pilot that can simultaneously address both ROI-level and WSI-level tasks across various anatomical regions while demonstrating robust pathological reasoning capability. b. We curate a dataset comprising 2,295,095 ROI-level and 188,154 WSI-level samples, enabling comprehensive model training and evaluation. c. Performance evaluation of MLLMs across 46 ROI-level tasks and 26 WSI-level tasks. * represents external validation tasks. SmartPath-R1 demonstrates exceptional performance over state-of-the-art MLLMs. d. The framework, comprising scale-dependent supervised fine-tuning and task-aware reinforcement fine-tuning, enables multimodal pathological reasoning within <think></think> tags. ROIs and WSIs are encoded into variable-length visual token sequences. Mixture-of-expert mechanism is used to address a spectrum of tasks, from ROI-level classification to WSI-level VQA.
@@
The emergence of multimodal large language models (MLLMs) [20, 21, 22, 23] has created new opportunities, yet two fundamental limitations persist. First, existing models often fail to achieve clinically meaningful reasoning. This is largely due to their reliance on supervised fine-tuning  [24, 22, 23], which optimizes the accuracy of prediction but fails to capture the sequential decision-making process inherent in pathological diagnosis, where experts progressively integrate visual cues and knowledge. As a result, models may learn superficial correlations rather than the critical reasoning pathways needed for complex cases. Second, existing approaches exhibit limited versatility in handling various computational pathology tasks. Although current MLLMs demonstrate preliminary capabilities in ROI-level VQA [22, 23], their capacity to concurrently process multiscale queries (ROI and WSI level) remains underdeveloped. These methods lack unified frameworks to integrate fundamental pathological tasks, including fine-grained classification, detection, and segmentation, with diagnostic reasoning. This fragmentation hinders the development of comprehensive diagnostic assistants capable of mirroring the multitask analytical workflows of pathologists, where visual interpretation and semantic reasoning are intrinsically coupled across spatial hierarchies.
@@
To overcome these limitations, we present SmartPath-R1 (Figure 1), an MLLM trained on extensive datasets comprising 1,964,121 ROI samples and 179,897 WSI samples, with a rigorous evaluation performed on 330,974 ROI and 8,257 WSI test samples. Our approach uniquely embeds diagnostic reasoning by formulating pathology interpretation as a sequential policy optimization process, where the model learns optimal evidence-gathering policies through reward signals mimicking pathologist decision trajectories. Crucially, our approach eliminates the need for expensive chain-of-thought supervision by mining the knowledge inherent in MLLMs, thereby enhancing scalability. Moreover, SmartPath-R1 unifies multiscale and multitask analysis via a mixture-of-expert mechanism that dynamically handles ROI-level morphometrics to WSI-level topological features for different tasks.
@@
Figure 2: Comprehensive Comparison of MLLMs across 72 Tasks. a. Task types evaluated by different MLLMs. CLS, DET, SEG, and VQA represent classification, detection, segmentation, and visual question answering, respectively. b. Average rank of MLLMs across 72 downstream tasks. The solid line and the dashed line represent the median and the mean, respectively. c. Ranking order of MLLMs across 46 ROI-level tasks. d. Ranking order of MLLMs on 26 WSI-level tasks. * represents external validation datasets. If a model achieves the best performance, its rank value is set to 1. Models with identical metric values receive the same rank (tied ranking).
@@ 2 Results
We evaluate various MLLMs on 72 tasks, encompassing 8 ROI-level classification tasks, 14 ROI-level pathological detection tasks, 14 ROI-level pathological segmentation tasks, 10 ROI-level VQA tasks, 9 WSI-level classification tasks, and 17 WSI-level VQA tasks. Since the tasks involved different types of evaluation metrics, we evaluate the overall performance of the MLLMs using an average ranking approach. The model with the best performance was ranked 1st, while the model with the lowest performance was ranked 5th. As shown in Figure 2, the SmartPath-R1 model achieves the top average rank score of 1.1 (ranked first in 68 tasks), outperforming the second-best model, PathoR1, which has a ranking score of 2.9 (ranked first in 3 tasks). The results clearly indicate that SmartPath-R1 achieves state-of-the-art performance and is much more generalizable compared to the other MLLMs.
@@
Pathological ROI-level classification is a fundamental task in computational pathology, aiming to automatically categorize pathology ROI images into predefined classes. Unlike generic image classification, this task demands strong reasoning capabilities to decipher subtle morphological patterns (e.g., nuclear pleomorphism, tissue architecture) and integrate domain-specific knowledge (e.g., grading criteria, staining artifacts). Although traditional deep learning approaches excel at pattern recognition, integrating explicit reasoning mechanisms is increasingly vital to bridge the gap between AI and clinical trustworthiness.
@@
In our comprehensive evaluation, SmartPath-R1 is benchmarked against four state-of-the-art MLLMs: Qwen2.5VL [25] (general purpose vision-language model), LLaVA-Med [26] (biomedical-focused MLLM), Quilt-LLaVA [23] and PathoR1 [27] (specialized for pathology). As shown in Figure 3a-i, the evaluation of MLLMs across 8 distinct tasks, including 2 external validation datasets, reveals important insights into their capabilities. Qwen2.5VL surprisingly outperforms several domain-specialized models across multiple pathological datasets. This finding can be attributed to its extensive pretraining on diverse vision-language tasks, which fosters more robust feature representations capable of generalizing to unseen pathological patterns. PathoR1’s relatively limited performance highlights the challenges of overspecialization in pathological AI systems. While designed specifically for ROI-level VQA, its constrained training scope and modest sample size render it less adaptable to broader diagnostic tasks. In contrast, SmartPath-R1 achieves particularly strong results, where it outperforms the second-best approach by 38.3% in average accuracy. This superior performance can be attributed to the models’ ability to capture tissue patterns and contextual relationships in these complex classification tasks.
@@
We present an example of classification task in Figure 3j. General purpose Qwen2.5VL struggles with domain-specific reasoning, as it prioritizes broad visual-language alignment over histopathologic logic. LLaVA-Med’s correct but terse output (“shows a tumor”) lacks explanatory depth, limiting its utility in clinical settings where justification is paramount. Quilt-LLaVA and PathoR1’s misclassification as (G) and (B) likely stems from isolated feature recognition without integrating stromal or nuclear atypia cues. SmartPath-R1 identifies the tissue as (J) Tumor, reasoning that the glandular morphology indicates advanced gastric adenocarcinoma, excluding early-stage or benign features (e.g., ulceration or regression tissue). This aligns with the clinical context of adenocarcinoma, where malignant glandular presence is a hallmark.
@@
Figure 3: Performance of MLLMs on ROI-Level Classification Tasks. a. Average performance of MLLMs based on accuracy (ACC) and F1 across 8 ROI-level classification tasks. b-i. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test [28]. Additional results are shown in Extended Data Table 3 and Tables 4-11. j. An example of ROI-level classification along with the results generated by various MLLMs.
@@
Figure 4: Performance of MLLMs on ROI-Level Detection Tasks. a. Average performance of MLLMs based on average precision at IoU threshold 0.5 (AP50) across 14 ROI-level detection tasks. b-o. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 13, Tables 14-27, Figure 9 and Figure 10. p. An example of ROI-level detection task along with the results generated by various MLLMs. Red rectangles indicate ground truth annotations, while predicted bounding boxes of each method are shown in their corresponding colors. The AP50 scores for each method are displayed in the lower-right corner, quantitatively demonstrating their detection accuracy relative to the ground truth.
@@
ROI-level detection involves the localization and classification of pathological entities (e.g., cells, nuclei, tumor regions) in images. This task serves as the foundation for automated diagnosis, prognostic assessment, and biomedical research, enabling high-throughput analysis of morphological features that are often imperceptible to human observers. The clinical significance of this task necessitates not only detection accuracy but also interpretive reasoning—the ability to contextualize morphological patterns within pathological paradigms. For instance, distinguishing malignant nuclei from benign counterparts requires reasoning about spatial relationships (e.g., tumor infiltration), texture heterogeneity, and diagnostic criteria. Traditional deep learning approaches, while proficient in pattern recognition, often lack explicit reasoning mechanisms.
@@
Figure 4 presents a comparative evaluation of SmartPath-R1 with existing MLLMs and a state-of-the-art medical image detection model BiomedParse. The analysis leverages 14 benchmark datasets spanning nuclear, cell, and tissue detection, focusing on Average Precision at an IoU threshold of 0.5 (AP50) as the primary metric. The results demonstrate that SmartPath-R1 achieves a significantly higher average AP50 across all datasets (Figure 4a), indicating superior detection performance. Dataset-specific performances (Figures 4b-o) further reveal that SmartPath-R1 exhibits particularly strong performance on all the datasets (p < 0.001). This suggests that SmartPath-R1 possesses a greater capacity for reasoning about the complex visual cues necessary for accurate detection.
@@
The qualitative examples are shown in Figure 4p. SmartPath-R1’s bounding boxes tightly enclose the glands, demonstrating a precise understanding of gland boundaries and morphology. In contrast, other models often produce inaccurate or incomplete bounding boxes, indicating a failure to reason about the subtle visual cues that distinguish glands from surrounding tissue. These findings suggest that SmartPath-R1’s enhanced reasoning capabilities contribute to its superior performance on ROI-level detection tasks, enabling it to accurately localize objects even in challenging histological contexts.
@@
Figure 5: Performance of MLLMs on ROI-Level Segmentation Tasks. For MLLMS, masks are generated by using the predicted objects’ coordinates as prompts for the pre-trained MedSAM model. a. Average performance of MLLMs based on Dice score across 14 ROI-level segmentation tasks. b-o. Model performance on specific tasks. * represents external validation datasets. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 28 and Tables 29-30. p. An example of ROI-level segmentation task, along with the results generated by various MLLMs. Red masks indicate ground truth annotations, while predicted masks of each method are shown in their corresponding colors. The Dice scores for each method are displayed in the lower-right corner, quantitatively demonstrating their segmentation accuracy relative to the ground truth.
@@
To assess the segmentation capabilities of SmartPath-R1, we conduct a comparative evaluation against existing models on 14 ROI-level segmentation datasets (Figure 5). For segmentation using MLLMs, the predicted objects’ coordinates are used as prompts for the pre-trained MedSAM [29] model to generate masks. The primary evaluation metric is the Dice score, which quantifies the overlap between predicted and ground truth segmentation masks. The results demonstrate that SmartPath-R1 achieves a significantly higher average Dice score across all datasets (Figures 5a) compared to other models, indicating superior segmentation performance. Analysis of dataset-specific performance (Figures 5b-o) reveals that SmartPath-R1 exhibits particularly strong performance on datasets such as BCSS, CoNIC2022, CRAG, DigestPath2019 Cell, DigestPath2019 Tissue, MIDOG Breast, MIDOG Neuroendocrine, Rings, SICAPv2, and WSS4LUAD (p < 0.001). This suggests that SmartPath-R1 possesses a greater capacity for reasoning about the complex visual cues necessary for accurate segmentation.
Figure 5p presents the qualitative examples. SmartPath-R1’s segmentation masks closely align with the lymphocyte boundaries, demonstrating a precise understanding of cell morphology. In contrast, other models often produce inaccurate or incomplete masks, indicating a failure to reason about the subtle visual cues that distinguish lymphocytes from surrounding tissue. These findings suggest that SmartPath-R1’s enhanced reasoning capabilities contribute to its superior performance on ROI-level segmentation tasks, enabling it to accurately delineate histological structures even in complex pathology images.
@@
Figure 6: Performance of MLLMs on ROI-Level VQA Tasks. a. Average performance of MLLMs based on accuracy (ACC) across 10 ROI VQA tasks. b-h. Model performance on specific tasks. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 31 and Table 32. i. An example of ROI-level VQA task along with the results generated by various MLLMs.
Pathological VQA is an advanced multimodal reasoning task that requires models to interpret medical images and generate accurate answers to free-form clinical questions. Unlike classification operates within a static label space (e.g., binary tumor categorization) per dataset, VQA dynamically adapts to diverse diagnostic inquiries, thereby emulating the flexible and context-driven nature of real-world clinical decision-making. This task not only demands fine-grained visual-textual alignment but also produces more interpretable and clinically actionable outputs, making it suited for complex diagnostic scenarios where reasoning and justification are critical.
To benchmark the performance of SmartPath-R1 against existing MLLMs on the fundamental ROI-level VQA task, we conduct a comparative evaluation on 10 datasets. Results demonstrate that SmartPath-R1 achieves a higher average accuracy (ACC = 0.738) compared to other models, including Qwen2.5VL (ACC = 0.450), LLaVA-Med (ACC = 0.272), Quilt-LLaVA (ACC = 0.339) and PathoR1 (ACC = 0.618) (Figures 6b). This suggests an advantage in general ROI-level understanding. Detailed analysis of task-specific performance (Figures 6b-h) reveals that SmartPath-R1 exhibits statistically significant (p < 0.05) improvements on tasks requiring more complex reasoning about histological structures.
In the illustrative ROI-level VQA example (Figure 6i), SmartPath-R1 correctly identifies the tissue type while other MLLMs fail, qualitatively demonstrating its enhanced reasoning capabilities. We attribute these gains to SmartPath-R1’s training on a large-scale dataset without procedural supervision, enabling the model to learn more robust features and develop more sophisticated reasoning skills.
@@
Figure 7 summarizes the accuracy of five MLLMs across 9 WSI classification tasks, including 3 external ones. SmartPath-R1 consistently achieves the highest average accuracy (15.5 % absolute margin over the second-best method; p < 0.001, Wilcoxon signed-rank test), and demonstrates superior robustness on both internal (TCGA-BRCA, TCGA-COAD, TCGA-HNSC, TCGA-LGG, TCGA-LUAD, TCGA-LUSC) and external validation sets (BRACS, CAMELYON, CPTAC-NSCLC).
In the glioma subtyping example (Figure 7f ), competing models falter due to either insufficient pathological knowledge or perceptual errors in feature recognition, leading to misclassifications. SmartPath-R1 instead performs a deliberate, step-wise exclusion: absence of mitoses, microvascular proliferation, and necrosis rules out Glioblastoma Multiforme; moderate hypercellularity and mild nuclear pleomorphism align with anaplastic astrocytoma. The model explicitly states “no evidence of… distinguishing it from glioblastoma multiforme”, yielding a coherent rationale that leads to the correct answer (B) Anaplastic Astrocytoma.
Figure 7: Performance of MLLMs on WSI-Level Classification Tasks. a. Average performance of MLLMs based on accuracy (ACC) and F1 across 9 WSI-level classification tasks. b-e. Model performance on specific tasks. * represents external validation datasets. Error bars represent 95% CI. The box limits represent the standard error. P-values are computed using a Wilcoxon signed-rank two-sided test. Additional results are shown in Extended Data Table 34 and Table 35. f. An example of WSI-level classification task along with the results generated by various MLLMs.